OpenStudy (anonymous):

use the laws of logarithms to solve the equation log(base 2)*(2x+5)=4

jimthompson5910 (jim_thompson5910):

\[\Large \log_{2}(2x+5)=4\] \[\Large 2x+5=2^4\] \[\Large 2x+5=16\] \[\Large 2x=16-5\] \[\Large 2x=11\] \[\Large x=\frac{11}{2}\]